My son is seven times older than my grandson and I am 12 times older than my grandson. If you add all our ages together, the sum is 100 years. Can you tell me my age in years?
Answer by stanbon(75887)

My son is seven times older than my grandson and I am 12 times older than my grandson. If you add all our ages together, the sum is 100 years. Can you tell me my age in years?
----------------------
Equations:
s = 7g
f = 12g
f + s + g = 100
-------
Substitute for s and for f and solve for ":
-------
12g + 7g + g = 100
-----
20g = 100
g = 5 (grandson's age now)
s = 7g = 35 (son's age now)
f = 12g = 60 (father's age now)
--------------------------------------
Cheers,
Stan H.
=================